OpenStudy (anonymous):

What is the value of x? If necessary, round your answer to two decimal places.

OpenStudy (anonymous):

What is the equation?

OpenStudy (hesan):

we use Pythagorean theorem here i.e H^2 = P^2 + B^2 here from the figure. B = base = 13 H = Hypotenuse = 26 P = ? as we put in the equation above... (26)^2 = P^2 + (13)^2 676 = P^2 + 169 676 - 169 = P^2 P^2 = 507 taking root P = 22.51 i.e P = 26

OpenStudy (hesan):

@flashb407 This is the whole solution possible... Are you getting it???

OpenStudy (anonymous):

@Hesan thanks but i got 22.52....

OpenStudy (hesan):

sorry it is 22.52...I have committed a mistake in rounding off...

OpenStudy (hesan):

Got it Now??